OpenStudy (anonymous):

systems with 3 variables, solve by substitution: x+5y-4z=-10 2x-y+5z=-9 2x-10y-5z=0 there is answer choices as well, I have been trying to work this out for about 2 hours now and I cannot think of anything else to try! please help!!!!

OpenStudy (anonymous):

hi insanity cat :)

OpenStudy (anonymous):

srry bethany

OpenStudy (anonymous):

you can use substitution to find the value of one variable we can find the value of x in the first variable easily x=-10+4z-5y and hello there

OpenStudy (anonymous):

haha its okay!

OpenStudy (anonymous):

bai i dont know to solve those problems....

OpenStudy (anonymous):

im still confused!! :/ the steps that I have were making since but it just never gave me the x, y, and z

OpenStudy (anonymous):

so lets substitute x 2x-y+5z=-9 2x-10y-5z=0 2(-10+4z-5y)-y+5z=-9 2(-10+4z-5y)-10y-5z=0 -20+8z-10y-y+5z=-9 -20+8z-10y-10y-5z=0 -11y+13z-20=-9 -10y+3z-20=0 -11y+13z=11 -10y+3z=20

OpenStudy (anonymous):

so x would be 20? .... wait no, because that's not an answer choice. its either 5, or -5 for x

OpenStudy (anonymous):

but I don't see how I would get either one!

OpenStudy (anonymous):

well if you plug the equation into a calculator with the capabilities to solve a system like this you get x=-1 y=-1 and z=0

OpenStudy (anonymous):

sorry x=-5

OpenStudy (anonymous):

what kind of calculator are you using?? lol I need it!

OpenStudy (jhannybean):

insanity, for the second equation after substitution you have -10y-10y = -20y , not -10y

OpenStudy (anonymous):

I have a TI-nspire with me

OpenStudy (anonymous):

i was just going to check over taht

OpenStudy (anonymous):

that*

OpenStudy (jhannybean):

So bethany, have you followed up to the substitution part from @insanitycat 's method?

OpenStudy (anonymous):

2x-y+5z=-9 2x-10y-5z=0 2(-10+4z-5y)-y+5z=-9 2(-10+4z-5y)-10y-5z=0 -20+8z-10y-y+5z=-9 -20+8z-10y-10y-5z=0 -11y+13z-20=-9 -20y+3z-20=0 -11y+13z=11 -20y+3z=20

OpenStudy (jhannybean):

Good.

OpenStudy (anonymous):

so you substitute the first equation into the second for x, right?

OpenStudy (jhannybean):

for the second and third.

OpenStudy (anonymous):

no you can get x from the first equation and then substitute the x-value into the 2nd and third

OpenStudy (anonymous):

I still don't see where -5 would come in for x

OpenStudy (anonymous):

we will get to that

OpenStudy (anonymous):

first we need to find y and z

OpenStudy (anonymous):

this uses the elimination method

OpenStudy (anonymous):

this kind of substitution would require elimination though in order to make it a two variable equation

OpenStudy (anonymous):

no it is supposed to be elimination

OpenStudy (jhannybean):

Ah... at top you stated substitution.

OpenStudy (anonymous):

so we substitute y=-1 into one of the equations to find z -11y+13z=11 -11(-1)+13z=11 13z+11=11 13z=0 z=0

OpenStudy (anonymous):

I typed it wrong... but I was getting the part where it becomes 2 problems wrong so you can solve it as two equations!

OpenStudy (jhannybean):

There you go, now you've got y=-1 and z = 0, do you understand how you achieved both of these values?

OpenStudy (anonymous):

the next prob is substitution, this one is elimination! sorry yall

OpenStudy (anonymous):

then we can substitute both the y and z values into the equation for x now x=-10+4z-5y x=-10+4(0)-5(-1) x=-10+5 x=-5

OpenStudy (jhannybean):

Thank you @insanitycat and @bethanyquintero good luck!

OpenStudy (anonymous):

ok wait, I understand everything on how to work the problem, but how you got -1 for x...

OpenStudy (jhannybean):

x = -5, y=-1.

OpenStudy (anonymous):

x=-5

OpenStudy (anonymous):

i see it now, so i see how to get the x and z but how did you get the y=-1..... if i can get that then i will completely understand the problem!